OpenStudy (yuii):

The ordered pairs model an exponential function, where j is the function name and e is the input variable. {(1, 10),  (2, 50) , ( 3, 250) , (4, 1250)} What is the function equation in sequence notation?

OpenStudy (welshfella):

still here? I thought Id helped you with this already!!

OpenStudy (yuii):

I didn't know.

OpenStudy (welshfella):

10(5)^(1-1) = 1 10(5)^(2-1) = 50 10(5)^(3-1)= 250 10(5)^(4-1)= 1250 so J = 10(5)^(e-1)

OpenStudy (emogirlatemoooocow):

@Yuii they Did this for you :/